# Biquadratic SOS Rank and Double Zarankiewicz Number

### 05 Mar 2026 v1

@@ Abstract
The maximum sum-of-squares (SOS) rank of an $m\times n$ biquadratic form, denoted $\operatorname{BSR}(m,n)$, has recently been shown to be bounded below by the classical Zarankiewicz number $z(m,n)$, the maximum number of edges in a $C_{4}$-free bipartite graph. However, a gap was discovered for the $4\times 3$ case: $\operatorname{BSR}(4,3)\geq 8>z(4,3)=7$, arising from a construction that adds a square of a bilinear form $(x_{4}y_{2}+x_{1}y_{3})^{2}$ to a $C_{4}$-free simple biquadratic form. This paper introduces the double Zarankiewicz number $z_{2}(m,n)$ to capture such phenomena. We define a new class of bipartite graphs that may contain both ordinary edges (1-edges) and “double edges” (2-edges), the latter corresponding to squares of simple two-term bilinear forms. A graph $G$ of this type yields a doubly simple biquadratic form $P_{G}$, and we prove that if $G$ contains no generalised $C_{4}$-cycle then $\operatorname{sos}(P_{G})=|E_{1}|+|E_{2}|$; consequently $\operatorname{BSR}(m,n)\geq z_{2}(m,n)$. We determine the exact value of $z_{2}(m,n)$ for several small parameters: $z_{2}(m,2)=m+1$, $z_{2}(2,n)=n+1$, $z_{2}(3,3)=6$, $z_{2}(4,3)=8$, and $z_{2}(5,3)=9$. For the $4\times 4$ case we establish $10\leq z_{2}(4,4)\leq 11$, leaving the exact value as an open problem. These results improve the known lower bounds for $\operatorname{BSR}(4,4)$ and $\operatorname{BSR}(5,3)$. The paper concludes with a discussion of open questions and connections between extremal graph theory and the study of SOS representations.
@@ 1 Introduction
If $G$ has no $C_{4}$-cycles, then the simple biquadratic form $P_{G}$ is irreducible, i.e., $\mathrm{sos}(P)=|E|$. On the other hand, the Zarankiewicz number $z(m,n)$ is defined as the largest number of edges such that the $m\times n$ bipartite graph $G$ is $C_{4}$-cycle free [10, 8]. Thus, we have (4).
@@
The remainder of this paper is organized as follows. In Section 2, we introduce the concept of a double Zarankiewicz number $z_{2}(m,n)$ and establish its basic properties, including its relation to the classical Zarankiewicz number and exact values for small parameters. We also prove the fundamental inequality $\operatorname{BSR}(m,n)\geq z_{2}(m,n)$ by showing that any graph attaining the double Zarankiewicz number yields an irreducible doubly simple biquadratic form. Section 3 investigates the $4\times 4$ case, presenting a construction achieving ten edges and establishing the bounds $10\leq z_{2}(4,4)\leq 11$. Section 4 treats the $5\times 3$ case, where we prove $z_{2}(5,3)=9$. Finally, Section 5 concludes the paper with a summary of results and a discussion of open problems.
@@ 2 Double Zarankiewicz Number
Let $G=(S,T,E)$ be an $m\times n$ bipartite graph, where $S=\{1,\cdots,m\}$ and $T=\{1,\cdots,n\}$ are its vertex sets. We define the edge set $E=E_{1}\cup E_{2}$, where $E_{1}$ is the $1$-edge set, and $E_{2}$ is the $2$-edge set. A $1$-edge $e$ in $E_{1}$ has the form $e=(i,j)$, where $i\in S$ and $j\in T$. On the other hand, a $2$-edge $e$ in $E_{2}$ has the form $(i,j;k,l)$, where $i,k\in S$ and $j,l\in T$. It is fine that $i=k$ or $j=l$ but should not be both $i=k$ and $j=l$. For such a bipartite graph $G$, we associate it with an $m\times n$ sos biquadratic form $P_{G}$, defined as:
@@
Let $m=4$ and $n=3$, $E_{1}=\{(1,1),(2,2),(3,3),(1,2),(2,3),(3,1),(4,1)$ and $E_{2}=\{(i,j;k,l)\}$. Then we see if $(i,j;k,l)$ is one of the following:
(4,2;1,3),(4,3;1,2),(4,2;4,3),(4,1;4,2),(4,2;1,3),\;(4,3;1,2),\;(4,2;4,3),\;(4,1;4,2),
$P_{G}$ is irreducible. We wish to characterize the combinatoric feature of $(i,j;k,l)$.
We now refine the notion of a $C_{4}$-cycle to account for the presence of 2-edges and formulate a linear independence condition that characterizes irreducibility.
@@
  • $E_{2}$ is a set of 2-edges of the form $(i,j;k,l)$ with $i,k\in S$ and $j,l\in T$ (allowing $i=k$ or $j=l$, but not both simultaneously).
Assume that there are no repeated 1-edges or 2-edges. We say that $G$ contains a generalized $C_{4}$-cycle if either
  1. (i)
    there exist distinct rows $i,k\in S$ and distinct columns $j,l\in T$ such that, counting each 1-edge as one ordinary edge and each 2-edge $(p,q;r,s)$ as two ordinary edges $(p,q)$ and $(r,s)$, the total number of ordinary edges (with multiplicities) contained in the set $\{i,k\}\times\{j,l\}$ is at least four (if the vertices of a 2-edge are not both in the block $\{i,k\}\times\{j,l\}$, then this 2-edge is not counted);
  2. (ii)
    there exists a nontrivial linear combination
    $\sum_{(p,q)\in E_{1}}\alpha_{pq}\,\mathbf{e}_{pq}\;+\!\!\sum_{(i,j;k,l)\in E_{2}}\beta_{ijkl}\,(\mathbf{e}_{ij}+\mathbf{e}_{kl})=\mathbf{0}$
    in $\mathbb{R}^{mn}$ (where $\mathbf{e}_{pq}$ is the standard basis vector corresponding to $x_{p}y_{q}$), with coefficients not all zero, and such that there exist distinct rows $r,s\in[m]$ and distinct columns $t,u\in[n]$ for which every basis vector $\mathbf{e}_{pq}$ appearing in the combination (i.e., with $\alpha_{pq}\neq 0$ or as part of a 2-edge term with $\beta_{ijkl}\neq 0$) satisfies $p\in\{r,s\}$ and $q\in\{t,u\}$.
@@
By definition, taking $|E_{2}|=0$ we have $z_{2}(m,n)\geq z(m,n)$ for all $m,n\geq 2$.
Now consider $(m,2)$. Let the two columns be $1$ and $2$. For any admissible graph $G=(S,T,E_{1}\cup E_{2})$ with $|T|=2$, define for each row $i\in S$ the total multiplicity $s_{i}$ of ordinary edges incident to that row, i.e., the number of times the cell $(i,1)$ appears plus the number for $(i,2)$, counting each 1-edge as one and each 2-edge $(p,q;r,s)$ as one ordinary edge for each of its two constituent cells. Then for any two distinct rows $i,k$, the submatrix $\{i,k\}\times\{1,2\}$ contains exactly $s_{i}+s_{k}$ ordinary edges. Condition (i) of Definition 2.1 requires $s_{i}+s_{k}\leq 3$ for every pair $i\neq k$; otherwise a $C_{4}$ would appear with at least four edges.
We now bound the total sum $M=\sum_{i=1}^{m}s_{i}$. If the largest $s_{i}$ equals $2$, then all other rows must have $s_{i}\leq 1$ (since $2+2=4$ is forbidden). Hence $M\leq 2+(m-1)\cdot 1=m+1$. If the largest equals $3$, then every other row must have $s_{i}=0$ (because $3+1=4$ would violate the condition), so $M\leq 3$. If all $s_{i}\leq 1$, then $M\leq m$. In all cases $M\leq m+1$.
The total number of edges $T=|E_{1}|+|E_{2}|$ satisfies $M=|E_{1}|+2|E_{2}|$ because each 2-edge contributes two ordinary edges. Hence $T=M-|E_{2}|\leq M\leq m+1$. Thus $z_{2}(m,2)\leq m+1$.
Conversely, the star graph where one row has both columns (i.e., two 1-edges) and every other row has a single 1-edge to column $1$ (or symmetrically) contains exactly $m+1$ edges and no $C_{4}$ (any two rows share at most one column). Hence it is admissible, giving $z_{2}(m,2)\geq m+1$. Therefore $z_{2}(m,2)=m+1$. By symmetry, interchanging rows and columns yields $z_{2}(2,n)=n+1$.
Now consider $(3,3)$. For a graph $G$ on $3$ rows and $3$ columns, there are $\binom{3}{2}^{2}=9$ choices of two distinct rows and two distinct columns. For each such $2\times 2$ submatrix, condition (i) limits the total number of ordinary edges (with multiplicities) contained in it to at most $3$. Summing over all nine submatrices we obtain a total $S\leq 9\times 3=27$.
Every ordinary edge $(i,j)$ (counted with its multiplicity) appears in exactly (31)(31)=4(3-1)(3-1)=4 of these submatrices (choose one other row and one other column). If $M$ denotes the total number of ordinary edges (i.e., $M=|E_{1}|+2|E_{2}|$), then $4M=S\leq 27$, so $M\leq 6.75$ and consequently $M\leq 6$. Since $T=|E_{1}|+|E_{2}|\leq M$, we obtain $T\leq 6$. Thus any admissible graph has at most $6$ edges. The Zarankiewicz number $z(3,3)=6$ is attained by a $C_{4}$-free bipartite graph (e.g., a $3\times 3$ grid with six edges); taking $E_{2}=\emptyset$ gives an admissible configuration with $6$ edges. Hence $z_{2}(3,3)=6$.
Finally, for $(4,3)$, we exhibit a construction with total edges $8$. Let
@@
Condition (i). For each pair of distinct rows $i,k\in\{1,2,3,4\}$ and distinct columns $j,l\in\{1,2,3\}$, we count the number of 1?edges in $\{i,k\}\times\{j,l\}$ plus twice the number of 2-edges whose two constituent ordinary edges both lie in this submatrix. The only 2-edge is $(4,2;1,3)$; its two ordinary edges are $(4,2)$ and $(1,3)$. They lie together in the submatrix $\{1,4\}\times\{2,3\}$. Hence in that submatrix we have one 1-edge $(1,2)$ (since $(1,2)\in E_{1}$) and one 2-edge, giving a total of $1+2=3$. In all other $2\times 2$ submatrices, no 2-edge is fully contained, and one checks directly that the number of 1-edges never exceeds $3$. For instance:
$\{i,k\}\times\{j,l\}$ 1-edges in submatrix
$\{1,2\}\times\{1,2\}$ $(1,1),(1,2),(2,2)$ (3)
$\{1,2\}\times\{2,3\}$ $(1,2),(2,2),(2,3)$ (3)
$\{1,3\}\times\{1,3\}$ $(1,1),(3,1),(3,3)$ (3)
$\{1,4\}\times\{1,2\}$ $(1,1),(1,2),(4,1)$ (3)
A complete verification shows every $2\times 2$ submatrix contains at most three ordinary edges (counting 2-edges only when fully present). Thus condition (i) does not hold.
Condition (ii). Consider any pair of distinct rows $i,k$ and distinct columns $j,l$. The vectors in $\mathbb{R}^{12}$ (with basis $\mathbf{e}_{pq}$) that arise from $E_{1}$ in this submatrix are the basis vectors $\mathbf{e}_{pq}$ for each $(p,q)\in E_{1}\cap\{i,k\}\times\{j,l\}$. Any 2-edge $(p,q;r,s)$ that is fully contained contributes the vector $\mathbf{e}_{pq}+\mathbf{e}_{rs}$. In our graph, the only 2?edge is $(4,2;1,3)$, which is fully contained only in the submatrix $\{1,4\}\times\{2,3\}$. There we have one 1-edge $(1,2)$ giving $\mathbf{e}_{12}$, and the 2-edge giving $\mathbf{e}_{13}+\mathbf{e}_{42}$. These two vectors are linearly independent because they involve distinct basis vectors. In every other submatrix, only 1-edges appear, and their corresponding basis vectors are linearly independent. Hence no nontrivial linear combination supported in any $\{i,k\}\times\{j,l\}$ can vanish. Therefore condition (ii) fails. Thus, the graph contains no generalized $C_{4}$-cycle. With $|E_{1}|+|E_{2}|=7+1=8$, we obtain $z_{2}(4,3)\geq 8$. As $z(4,3)=7$, we have $z_{2}(4,3)>z(4,3)$. ∎

### 16 Mar 2026 v2

@@ Abstract
The maximum sum-of-squares (SOS) rank of an $m\times n$ biquadratic form, denoted $\operatorname{BSR}(m,n)$, has recently been shown to be bounded below by the classical Zarankiewicz number $z(m,n)$, the maximum number of edges in a $C_{4}$-free bipartite graph. However, a gap was discovered for the $4\times 3$ case: $\operatorname{BSR}(4,3)\geq 8>z(4,3)=7$, arising from a construction that adds a square of a bilinear form $(x_{4}y_{2}+x_{1}y_{3})^{2}$ to a $C_{4}$-free simple biquadratic form. This paper introduces the double Zarankiewicz number $z_{2}(m,n)$ to capture such phenomena. We define a new class of bipartite graphs that may contain both ordinary edges (1-edges) and “double edges” (2-edges), the latter corresponding to squares of simple two-term bilinear forms. A graph $G$ of this type yields a doubly simple biquadratic form $P_{G}$, and we prove that if $G$ contains no generalised $C_{4}$-cycle then $\operatorname{sos}(P_{G})=|E_{1}|+|E_{2}|$; consequently $\operatorname{BSR}(m,n)\geq z_{2}(m,n)$. We determine the exact value of $z_{2}(m,n)$ for several small parameters: $z_{2}(m,2)=m+1$, $z_{2}(2,n)=n+1$, $z_{2}(3,3)=6$, $z_{2}(4,3)=8$, $z_{2}(4,4)=10$, and $z_{2}(5,3)=9$. Furthermore, we establish the lower bound $z_{2}(8,7)\geq 39$, which implies $\operatorname{BSR}(8,7)\geq 39$, improving the previous bound $\operatorname{BSR}(8,7)\geq z(8,7)=37$. These results demonstrate that the double Zarankiewicz number can exceed the classical one by at least two, revealing new algebraic-combinatorial phenomena. The paper concludes with a discussion of open questions and connections between extremal graph theory and the study of SOS representations.
@@ 1 Introduction
If $G$ has no $C_{4}$-cycles, then the simple biquadratic form $P_{G}$ is irreducible, i.e., $\mathrm{sos}(P_{G})=|E|$. On the other hand, the Zarankiewicz number $z(m,n)$ is defined as the largest number of edges such that the $m\times n$ bipartite graph $G$ is $C_{4}$-cycle free [13, 11]. Thus, we have (4).
@@
The remainder of this paper is organized as follows. In Section 2, we introduce the concept of a double Zarankiewicz number $z_{2}(m,n)$ and establish its basic properties, including its relation to the classical Zarankiewicz number and exact values for small parameters. We also prove the fundamental inequality $\operatorname{BSR}(m,n)\geq z_{2}(m,n)$ by showing that any graph attaining the double Zarankiewicz number yields an irreducible doubly simple biquadratic form. Section 3 investigates the $4\times 4$ case, where we prove $z_{2}(4,4)=10$. Section 4 treats the $5\times 3$ case, where we prove $z_{2}(5,3)=9$. Section 5 presents a new lower bound for $z_{2}(8,7)$, showing that $z_{2}(8,7)\geq 39=z(8,7)+2$ and consequently $\operatorname{BSR}(8,7)\geq 39$. Finally, Section 6 concludes the paper with a summary of results and a discussion of open problems.
@@ 2 Double Zarankiewicz Number
Let $G=(S,T,E)$ be an $m\times n$ bipartite graph, where $S=\{1,\cdots,m\}$ and $T=\{1,\cdots,n\}$ are its vertex sets. We define the edge set $E=E_{1}\cup E_{2}$, where $E_{1}$ is the $1$-edge set, and $E_{2}$ is the $2$-edge set. A $1$-edge $e$ in $E_{1}$ has the form $e=(i,j)$, where $i\in S$ and $j\in T$. On the other hand, a $2$-edge $e$ in $E_{2}$ has the form $(i,j;k,l)$, where $i,k\in S$ and $j,l\in T$. A $2$-edge is required to be nondegenerate, i.e., $i\neq k$ and $j\neq l$. For such a bipartite graph $G$, we associate it with an $m\times n$ sos biquadratic form $P_{G}$, defined as:
@@
Let $m=4$ and $n=3$, $E_{1}=\{(1,1),(2,2),(3,3),(1,2),(2,3),(3,1),(4,1)$ and $E_{2}=\{(i,j;k,l)\}$. Then we see if $(i,j;k,l)$ is $(4,2;1,3)$, $P_{G}$ is irreducible. We wish to characterize the combinatorial feature of such $(i,j;k,l)$.
We now refine the notion of a $C_{4}$-cycle to account for the presence of 2-edges.
@@
  • $E_{2}$ is a set of 2-edges of the form $(i,j;k,l)$ with $i,k\in S$ and $j,l\in T$. A 2-edge is required to be nondegenerate, i.e. $i\neq k$ and $j\neq l$.
We impose the following simplicity condition:
  1. (S)
    No two distinct edges or subedges of $G$ share a cell. We also assume that $|E_{1}|=z(m,n)$.
We say that $G$ contains a generalized $C_{4}$-cycle if either
  1. 1.
    there exists a classical $C_{4}$-cycle, or
  2. 2.
    there exists a 2-edge $(i,j;k,l)\in E_{2}$ such that both opposite cells $(i,l)$ and $(k,j)$ are occupied, or
  3. 3.
    there exist a 2-edge $(i,j;p,q)$ and a distinct cell $(k,l)$ such that the five cells $(k,l),(k,j),(k,q),(i,l),(p,l)$ are occupied.
@@
Remark 2.2.
Like condition 1, conditions 2 and 3 are also natural and combinatorial. Condition 2 ensures that a 2 edge itself does not create a $C_{4}$ cycle. Because $|E_{1}|=z(m,n)$, each half of a 2 edge would form a $C_{4}$ cycle with some distinct occupied cells (otherwise more 1-edges could be added). Condition?3 then guarantees that the two lists of such cells those completing a $C_{4}$ with the first half and those completing a $C_{4}$ with the second half have no intersection. This symmetric condition is exactly what prevents the linear dependence that would otherwise arise.
Theorem 2.3.
For all $m,n\geq 2$, we have $\operatorname{BSR}(m,n)\geq z_{2}(m,n)\geq z(m,n)$.
Proof.
The inequality $z_{2}(m,n)\geq z(m,n)$ is trivial by taking $E_{2}=\emptyset$. Now let $G=(S,T,E_{1}\cup E_{2})$ be an $m\times n$ bipartite graph attaining $z_{2}(m,n)$; thus $|E_{1}|+|E_{2}|=z_{2}(m,n)$ and $G$ contains no generalized $C_{4}$-cycle in the sense of Definition 2.1. Consider the associated doubly simple biquadratic form
Assume, for contradiction, that $\mathrm{sos}(P_{G})<|E_{1}|+|E_{2}|$. Then there exists an integer $r<|E_{1}|+|E_{2}|$ and bilinear forms
such that
For each pair $(i,j)$ define the vector $\mathbf{v}_{ij}=(a_{ij}^{(1)},\dots,a_{ij}^{(r)})\in\mathbb{R}^{r}$. If $(i,j)$ is not a 1-edge and not a half of any 2-edge, then comparing coefficients shows $\mathbf{v}_{ij}=\mathbf{0}$. Let
[…]
By the simplicity condition (S), the $2|E_{2}|$ halves are all distinct and none belongs to $E_{1}$; hence
$|\mathcal{V}|=|E_{1}|+2|E_{2}|\geq|E_{1}|+|E_{2}|>r.$
Thus the vectors in $\mathcal{V}$ are linearly dependent. Choose a nontrivial linear relation
$\sum_{(i,j)\in\mathcal{V}}\alpha_{ij}\mathbf{v}_{ij}=\mathbf{0}$
with minimal support $S=\{(i,j):\alpha_{ij}\neq 0\}$. All vectors in this relation are nonzero and, by minimality, they are distinct (if two were equal we could combine them and obtain a smaller support).
  • For every 1-edge $(i,j)\in E_{1}$: $\|\mathbf{v}_{ij}\|^{2}=1$. (A1)
  • For every 2-edge $(i,j;k,l)\in E_{2}$: $\|\mathbf{v}_{ij}\|^{2}=\|\mathbf{v}_{kl}\|^{2}=1$ and $\mathbf{v}_{ij}\!\cdot\!\mathbf{v}_{kl}+\mathbf{v}_{il}\!\cdot\!\mathbf{v}_{kj}=1$. (A2)
  • For any two distinct pairs $(p,q),(r,s)$ that are not the two halves of the same 2-edge: $\mathbf{v}_{pq}\!\cdot\!\mathbf{v}_{rs}+\mathbf{v}_{ps}\!\cdot\!\mathbf{v}_{rq}=0$. (B)
  • For pairs sharing a row or a column: $\mathbf{v}_{ij}\!\cdot\!\mathbf{v}_{il}=0$ (same row), $\mathbf{v}_{ij}\!\cdot\!\mathbf{v}_{kj}=0$ (same column). (C)
Step 1. Consequences of the absence of a generalized $C_{4}$-cycle. Take any 2-edge $(i,j;k,l)\in E_{2}$. Because it is nondegenerate, $i\neq k$ and $j\neq l$. Look at the submatrix $\{i,k\}\times\{j,l\}$. Since no generalized $C_{4}$-cycle exists, at most one of the two opposite cells $(i,l)$ and $(k,j)$ can be occupied (otherwise condition 2 of Definition 2.1 would be violated). Hence at least one of $\mathbf{v}_{il},\mathbf{v}_{kj}$ is zero. From (A2) we then obtain $\mathbf{v}_{ij}\!\cdot\!\mathbf{v}_{kl}=1$. Because $\|\mathbf{v}_{ij}\|=\|\mathbf{v}_{kl}\|=1$, Cauchy-Schwarz forces $\mathbf{v}_{ij}=\mathbf{v}_{kl}$. Thus each 2-edge contributes two equal vectors; denote this common vector by $\mathbf{w}_{ijkl}$.
Step 2. Orthogonality of distinct vectors. Let $(i,j)$ and $(k,l)$ be two distinct positions in the support $\mathcal{S}$. Apply relation (B) to these two positions:
$\mathbf{v}_{ij}\!\cdot\!\mathbf{v}_{kl}+\mathbf{v}_{il}\!\cdot\!\mathbf{v}_{kj}=0.$
Suppose, for contradiction, that $\mathbf{v}_{ij}\!\cdot\!\mathbf{v}_{kl}\neq 0$. Then $\mathbf{v}_{il}\!\cdot\!\mathbf{v}_{kj}\neq 0$, so both $\mathbf{v}_{il}$ and $\mathbf{v}_{kj}$ are nonzero. Hence the cells $(i,l)$ and $(k,j)$ are occupied (either as 1-edges or as halves of 2-edges). Consequently the four cells
(i,j),(k,l),(i,l),(k,j)(i,j),\;(k,l),\;(i,l),\;(k,j)
are all occupied.
If all four were 1-edges, they would form a classical $C_{4}$, violating condition 1 of Definition 2.1. Therefore at least one of them is a half of a 2-edge. Without loss of generality, assume $(i,j)$ is a half of a 2-edge $e=(i,j;p,q)$ with $p\neq i,\;q\neq j$. Then $\mathbf{v}_{ij}=\mathbf{v}_{pq}$ and $(p,q)$ is occupied.
Now consider the pair $(p,q)$ and $(k,l)$. Their dot product is $\mathbf{v}_{pq}\!\cdot\!\mathbf{v}_{kl}=\mathbf{v}_{ij}\!\cdot\!\mathbf{v}_{kl}\neq 0$. Apply (B) to $(p,q)$ and $(k,l)$:
$\mathbf{v}_{pq}\!\cdot\!\mathbf{v}_{kl}+\mathbf{v}_{pl}\!\cdot\!\mathbf{v}_{kq}=0,$
so $\mathbf{v}_{pl}\!\cdot\!\mathbf{v}_{kq}\neq 0$; thus $(p,l)$ and $(k,q)$ are occupied.
We examine the possibilities for $p$ and $q$ relative to $k$ and $l$.
  • If $p=k$, then $(p,l)=(k,l)$ and $(p,q)=(k,q)$. But $(k,q)$ and $(k,l)$ share the same row, so by (C) we have $\mathbf{v}_{kq}\!\cdot\!\mathbf{v}_{kl}=0$. Yet $\mathbf{v}_{kq}=\mathbf{v}_{pq}=\mathbf{v}_{ij}$ and $\mathbf{v}_{ij}\!\cdot\!\mathbf{v}_{kl}\neq 0$, a contradiction.
  • If $q=l$, then $(p,l)$ and $(k,l)$ share the same column, so (C) gives $\mathbf{v}_{pl}\!\cdot\!\mathbf{v}_{kl}=0$ while $\mathbf{v}_{pl}=\mathbf{v}_{ij}$ again contradicts $\mathbf{v}_{ij}\!\cdot\!\mathbf{v}_{kl}\neq 0$.
Hence we must have $p\neq k$ and $q\neq l$. In this situation the five cells
(k,l),(k,j),(k,q),(i,l),(p,l)(k,l),\;(k,j),\;(k,q),\;(i,l),\;(p,l)
are all distinct (because $i\neq k$, $j\neq l$, $p\neq k$, $q\neq l$). They are occupied: $(k,l)$ and $(i,l)$ from the original four, $(k,j)$ from the original, $(k,q)$ and $(p,l)$ from the new ones. Moreover, $(i,j;p,q)$ is a 2-edge. Thus the configuration satisfies condition 3 of Definition 2.1, which is forbidden. This contradiction shows that our assumption $\mathbf{v}_{ij}\!\cdot\!\mathbf{v}_{kl}\neq 0$ is impossible. Therefore
$\mathbf{v}_{ij}\!\cdot\!\mathbf{v}_{kl}=0\hskip 28.80008pt\text{for any two distinct positions }(i,j),(k,l)\in\mathcal{S}.$
Hence all distinct vectors in the support are pairwise orthogonal.
Step 3. Contradiction. The vectors in the support are nonzero (by (A1) and the fact that halves of 2-edges have norm 1) and pairwise orthogonal, so they are linearly independent. But they satisfy the nontrivial linear relation (1), which is impossible. This contradiction shows that our assumption $\mathrm{sos}(P_{G})<|E_{1}|+|E_{2}|$ was false. Hence $\mathrm{sos}(P_{G})=|E_{1}|+|E_{2}|=z_{2}(m,n)$, proving $\mathrm{BSR}(m,n)\geq z_{2}(m,n)$. ∎
@@
First, for $m=2$ or $n=2$, we have the classical Zarankiewicz numbers $z(m,2)=m+1$ and $z(2,n)=n+1$. By Theorem 2.3, $z_{2}(m,n)\geq z(m,n)$. On the other hand, it is known that $\operatorname{BSR}(m,2)=m+1$ and $\operatorname{BSR}(2,n)=n+1$ [1]. Applying Theorem 2.3 again gives $\operatorname{BSR}(m,n)\geq z_{2}(m,n)$; hence $z_{2}(m,2)\leq m+1$ and $z_{2}(2,n)\leq n+1$. Consequently $z_{2}(m,2)=m+1$ for all $m\geq 2$ and $z_{2}(2,n)=n+1$ for all $n\geq 2$.
For $(3,3)$, the Zarankiewicz number is $z(3,3)=6$ and it is known that $\operatorname{BSR}(3,3)=6$ [2]. The same argument yields $z_{2}(3,3)\geq 6$ and $z_{2}(3,3)\leq\operatorname{BSR}(3,3)=6$; therefore $z_{2}(3,3)=6$.
Finally, for $(4,3)$ we exhibit a construction with total edges $8$. Let
@@
Condition 1. The 1-edges alone form a $C_{4}$-free graph because they attain the Zarankiewicz number $z(4,3)=7$; any two rows share at most one column.
Condition 2. For the unique 2-edge $(4,2;1,3)$, its opposite cells are $(4,3)$ and $(1,2)$. Here $(1,2)\in E_{1}$ but $(4,3)$ is unoccupied, so not both opposite cells are occupied.
Condition 3. We must check that there is no cell $(k,l)$ with $k\in\{1,2,3,4\},\;l\in\{1,2,3\}$ such that the five cells
(k,l),(k,2),(k,3),(4,l),(1,l)(k,l),\;(k,2),\;(k,3),\;(4,l),\;(1,l)
are all occupied. Observe that $(k,2)$ and $(k,3)$ are both occupied only for $k=1$ and $k=2$ (since $(3,2)$ and $(4,3)$ are missing). - For $k=1$: $(1,l)$ is occupied for every $l$, but $(4,l)$ is occupied only for $l=1,2$. * $l=1$ gives the cells $(1,1),(1,2),(1,3),(4,1)$ - only four distinct cells. * $l=2$ gives $(1,2),(1,2),(1,3),(4,2)$ - again four distinct cells. - For $k=2$: $(2,l)$ is occupied for $l=2,3$ (since $(2,1)$ is empty). * $l=2$ gives $(2,2),(2,2),(2,3),(4,2),(1,2)$ - four distinct cells. * $l=3$ would require $(4,3)$, which is missing. Thus no choice yields five *distinct* occupied cells. (The definition of a generalized $C_{4}$-cycle implicitly requires the five cells to be distinct, as used in the proof of Theorem 2.3; otherwise the configuration would already be covered by simpler conditions.)
Therefore $G$ contains no generalized $C_{4}$-cycle. With $|E_{1}|+|E_{2}|=7+1=8$, we obtain $z_{2}(4,3)\geq 8$. Since $z(4,3)=7$, we have $z_{2}(4,3)>z(4,3)$.